Solve the system of equations by graphing. Then classify the system as consistant or inconsistant, dependent or indepenent?
9x-9y=63
9y-9x=-63
Help me please with this one.

Answer by MathTutorer(13)

This is a easy version of a simultaneous equation.
There are two ways to solve this kind of equation.
Substitution and Elimination
We will use substitution here.

9x-9y=63
9y-9x=-63
/ top equation by 9
x-y=7
x=y+7
then substitute into bottom equation.
9y-9(y+7)=-63
9y-9y-63=-63
0y=0
y is undefined.
x=7
hope this helps.